# Guiding Masked Representation Learning to Capture Spatio-Temporal Relationship of Electrocardiogram

### 02 Feb 2024 v1

@@ 7 Conclusion
In this work, we propose ST-MEM, Spatio-Temporal Masked Electrocardiogram Modeling, to learn the general ECG representation, generally applicable to diverse ECG problems by incorporating the spatial and temporal relationship of ECG signal. Through extensive experiments, we first demonstrate that ECG representation learned from ST-MEM can exhibit spatial relationships. Since limb leads are measured in the frontal plane, while precordial leads are in the horizontal plane, each representation from ST-MEM reflects that spatial information. Second, we observe the temporal relationship through an attention map. Third, we evaluate the general ECG representation in various experiments, such as reduced lead and low-resource settings. We believe that our work, providing general ECG representation encapsulating the spatio-temporal relationship, can benefit the recent healthcare industry.
8 Future work
We believe that our methodology can be applied to various multivariate time-series domains. An example of this adaptability is the transformation of lead embedding into a general time-series feature embedding. In Appendix J, we demonstrate an additional experiment utilizing the human activity recognition (HAR) dataset. As part of our future work, we aim to explore the application of ST-MEM to other multivariate time-series datasets. Moreover, we will explore the various transfer learning approaches (e.g., partially fine-tuning the pre-trained model (He et al., 2022)) to adapt to downstream tasks using general ECG representation instead of naively fine-tuning the model.
@@ 9 Reproducibility Statement
This section provides reviewers to reproduce our experiments.
In our code, we include the implementation of ST-MEM and some other baselines which are written in Python. Moreover, we provide the pre-trained model parameters such that the others can fine-tune the model and obtain reproducible results. We give only one dataset PTB-XL that is already preprocessed and divided into the train, valid, and test; however, we provide all links so that the reviewer can download the rest of the datasets. Lastly, we provide details regarding training hyperparameters and schemes with hardware information.
@@
In Secion 4.1, there is all the information of datasets including the number of ECGs, the number of patients, sample rate, and its duration. Moreover, all ECG data are resampled to 250 Hz, and additional information can be found in Appendix A. Moreover, one of the downstream datasets, PTB-XL which is separated into the train, valid, and test, can be downloaded from link in our attached codes (README.md).
Even only preprocessed PTB-XL is given, all raw data can be downloaded from the below link:
Some details are not described in this paper. Ningbo and Chapman can be downloaded together in PhysioNet Challenge 2021(45,152 ECG recordings), and for CODE-15, as we mentioned above, ECGs with a duration of less than 10 seconds should be dropped out, and there are some zero padding at the start and end of the recordings, these zero-paddings should be stripped out.
9.2 codes
Reviewers can run the models through the files we have attached. Please read ’README.md’ file in our attachment.
9.3 hyperparameters
All details in pre-training, fine-tuning, and linear evaluation can be found in Appendix E. It includes basic hyperparameters such as epochs, batch size, and optimizer for information on hardware.
@@ Appendix A Details in experimental settings
In this section, we explain the downstream datasets and the preprocessing for both pre-training and downstream datasets. The evaluation details of downstream datasets are explained.
A.1 Data preprocessing
@@
Resampling all ECG data into 250Hz. We have a total of 6 ECG datasets, but their sampling frequencies are not the same. Chapman, Ningbo, CPSC2018, and PTB-XL have a sampling frequency of 500Hz, CODE-15 has a frequency of 400Hz, and PhysioNet2017 has a frequency of 200Hz. To standardize the ECG datasets, we resampled them all to a common sampling rate of 250Hz.
@@
A.2 Evaluation on downstream tasks
@@
Appendix B Additional analysis of ECG representation learned from ST-MEM
Appendix C Additional ablation study
Table 12: Ablation on decoder depth.
Decoder depth AUROC
1 0.925
4 0.933
8 0.931
11 0.931
Figure 7: Ablation on masking ratio.
Table 13: Ablation on lead indicating modules.
Ablation AUROC
[SEP] embedding Lead embedding
0.822
0.830
0.820
0.838
Appendix D Additional analysis of self-attention maps of ST-MEM
Appendix E Implementation details
Table 14: Hyperparameter settings.
Pre-training Fine-tuning Linear evaluation
Backbone ViT-B ViT-B ViT-B
Learning rate 0.0012 0.001 0.001
Batch size 2048 1024 32
Epochs 800 100 100
Optimizer AdamW AdamW AdamW
Warump steps 40 5 5
Appendix F Experiments of different problem settings for downstream dataset PTB-XL
@@
Appendix G Selecting augmentation for pre-training MoCo v3
  1. 1.
    Erase: Randomly setting the values of a chosen lead to 0.
  2. 2.
    Flip: Randomly inverting the signal vertically.
  3. 3.
    Drop: Randomly zeroing out signal values.
  4. 4.
    Cutout: Selecting a random interval and setting its values to 0.
  5. 5.
    Shift: Randomly shifting the signal temporally.
  6. 6.
    Sine: Adding a sine wave to the entire signal.
  7. 7.
    Partial sine: Adding a sine wave into a randomly selected interval.
  8. 8.
Methods PTB-XL CPSC2018
Accuracy F1 AUROC Accuracy F1 AUROC
Linear Evaluation
MoCo v3, 4 augmentations 0.552 0.142 0.709 0.209 0.038 0.644
MoCo v3, 8 augmentations 0.552 0.142 0.739 0.268 0.08 0.712
Fine-tuning
MoCo v3, 4 augmentations 0.798 0.636 0.915 0.833 0.816 0.967
MoCo v3, 8 augmentations 0.799 0.644 0.910 0.852 0.838 0.967
Appendix H Additional single lead experiments on various lead types
@@
Appendix I Results of single lead ECGs in a low-resource setting
@@
Appendix J Applying ST-MEM to other multivariate time-series data
The Human Activity Recognition (HAR) dataset (Anguita et al., 2013) is a multivariate time series acquired from 30 individuals engaged in daily activities while wearing a waist-mounted smartphone equipped with inertial sensors. Data were collected in 2.56 seconds with a sampling frequency of 50Hz, resulting in 128 readings per sample. The collected data include triaxial acceleration from the accelerometer, estimated triaxial body acceleration, and triaxial angular velocity from the gyroscope, constituting a total of 9 features. The goal is to classify six different activities: walking, walking upstairs, walking downstairs, sitting, standing, and lying down.
@@
Appendix K Comparison of existing ECG pre-trained models with ST-MEM
Methods PTB-XL CPSC2018
1% 5% 100% 1% 5% 100%

### 11 Mar 2024 v2

@@ 7 Conclusion
In this work, we propose ST-MEM, Spatio-Temporal Masked Electrocardiogram Modeling, to learn the general ECG representation, generally applicable to diverse ECG problems by incorporating the spatial and temporal relationship of ECG signal. Through extensive experiments, we first demonstrate that ECG representation learned from ST-MEM can exhibit spatial relationships. Since limb leads are measured in the frontal plane, while precordial leads are in the horizontal plane, each representation from ST-MEM reflects that spatial information. Second, we observe the temporal relationship through an attention map. Third, we evaluate the general ECG representation in various experiments, such as reduced lead and low-resource settings. We believe that our work, providing general ECG representation encapsulating the spatio-temporal relationship, can benefit the recent healthcare industry. As part of our future work, we aim to explore the application of ST-MEM to general multivariate time-series datasets. Our experimental results on the human activity recogtion (Anguita et al., 2013) task can be found in Appendix C. Moreover, we will explore the various transfer learning approaches (e.g., partially fine-tuning the pre-trained model (He et al., 2022)) to adapt to downstream tasks using general ECG representation instead of naively fine-tuning the model.
@@ 8 Reproducibility Statement
Our code encompasses the implementation of ST-MEM and several other baselines written in Python. Furthermore, we furnish the pre-trained model parameters to facilitate others in fine-tuning the model and achieving reproducible results. Moreover, comprehensive details on training hyperparameters, schemes, and hardware specifications are provided. While we provide just one downstream dataset, PTB-XL, already preprocessed and partitioned into train, validation, and test sets, we ensure access to all additional datasets via provided links.
8.1 Codes
The codes will be available at https://github.com/bakqui/ST-MEM. Please read the ’README.md’ file and follow the instructions.
8.2 Hyperparameters
All details in pre-training, fine-tuning, and linear evaluation can be found in Appendix A.1. It includes basic hyperparameters such as epochs, batch size, and optimizer.
@@ 8.3 Datasets
In Secion 4.1, there is all the information of datasets including the number of ECGs, the number of patients, sample rate, and its duration.
Moreover, all ECG data are resampled to 250 Hz, and additional information can be found in Appendix A.2. Moreover, one of the downstream datasets, PTB-XL, separated into the train, valid, and test datasets, can be downloaded from the link in our codes (README.md).
Although only preprocessed PTB-XL is given, all raw data can be downloaded from the below links:
Ningbo and Chapman can be downloaded together in PhysioNet Challenge 2021 (45,152 ECG recordings). For CODE-15, as we mentioned in Secion 4.1, ECGs with a duration of less than 10 seconds should be dropped out, and there are some zero padding at the start and end of the recordings, these zero-paddings should be stripped out.
@@ Appendix A Details in experimental settings
In this section, we provide in-depth explanations of our experimental setup, including our approach to dataset processing, model evaluation on downstream tasks, and implementation of baseline methodologies.
A.1 Implementation details
Table 6: Hyperparameter settings.
Pre-training Fine-tuning Linear evaluation
Backbone ViT-B ViT-B ViT-B
Learning rate 0.0012 0.001 0.001
Batch size 2048 1024 32
Epochs 800 100 100
Optimizer AdamW AdamW AdamW
Warump steps 40 5 5
A.2 Data preprocessing
@@
ECG signal processing We have 6 ECG datasets, each with different characteristics. For instance, Chapman and CODE-15 have varying sampling frequencies (500Hz and 400Hz) and units (microvolts and millivolts). These differences in data resolution and scale could potentially confuse the model and diminish prediction performance. To ensure consistency among datasets and improve ECG signal quality, we implemented several signal processing procedures. Initially, we resampled all ECGs to a uniform sampling frequency of 250Hz. Next, we applied a bandpass digital filter with a range of 0.67-40Hz to eliminate baseline wandering and high-frequency noise. Finally, we normalized the ECG signals using Z-normalization to ensure uniform scales. The signal processing was conducted for all baselines regardless of whether it was during the pre-training or fine-tuning stage.
@@
A.3 Evaluation on downstream tasks
@@
Appendix B Additional experiments and results
B.1 Comparison of existing ECG pre-trained models with ST-MEM
Methods PTB-XL CPSC2018
1% 5% 100% 1% 5% 100%
B.2 Selecting augmentation for pre-training MoCo v3
  1. 1.
    Erase: Randomly setting the values of a chosen lead to 0.
  2. 2.
    Flip: Randomly inverting the signal vertically.
  3. 3.
    Drop: Randomly zeroing out signal values.
  4. 4.
    Cutout: Selecting a random interval and setting its values to 0.
  5. 5.
    Shift: Randomly shifting the signal temporally.
  6. 6.
    Sine: Adding a sine wave to the entire signal.
  7. 7.
    Partial sine: Adding a sine wave into a randomly selected interval.
  8. 8.
B.3 Experiments of different problem settings for downstream dataset PTB-XL
@@
Methods PTB-XL CPSC2018
Accuracy F1 AUROC Accuracy F1 AUROC
Linear Evaluation
MoCo v3, 4 augmentations 0.552 0.142 0.709 0.209 0.038 0.644
MoCo v3, 8 augmentations 0.552 0.142 0.739 0.268 0.08 0.712
Fine-tuning
MoCo v3, 4 augmentations 0.798 0.636 0.915 0.833 0.816 0.967
MoCo v3, 8 augmentations 0.799 0.644 0.910 0.852 0.838 0.967
@@
B.4 Additional single lead experiments on various lead types
@@
B.5 Results of single lead ECGs in a low-resource setting
@@
B.6 Additional analysis of ECG representation learned from ST-MEM
B.7 Additional ablation study
Table 20: Ablation on decoder depth.
Decoder depth AUROC
1 0.925
4 0.933
8 0.931
11 0.931
Figure 7: Ablation on masking ratio.
Table 21: Ablation on lead indicating modules.
Ablation AUROC
[SEP] embedding Lead embedding
0.822
0.830
0.820
0.838
B.8 Additional analysis of self-attention maps of ST-MEM
Appendix C Future work: applying ST-MEM to other multivariate time-series data
We believe that our methodology can be applied to various multivariate time-series domains. An example of this adaptability is the transformation of lead embedding into a general time-series feature embedding. In this section, we demonstrate an additional experiment utilizing the Human Activity Recognition (HAR) dataset (Anguita et al., 2013).
HAR dataset is a multivariate time series acquired from 30 individuals engaged in daily activities while wearing a waist-mounted smartphone equipped with inertial sensors. Data were collected in 2.56 seconds with a sampling frequency of 50Hz, resulting in 128 readings per sample. The collected data include triaxial acceleration from the accelerometer, estimated triaxial body acceleration, and triaxial angular velocity from the gyroscope, constituting a total of 9 features. The goal is to classify six different activities: walking, walking upstairs, walking downstairs, sitting, standing, and lying down.